# Identifying Causal Direction via Variational Bayesian Compression

### v3

@@ Appendix F Detailed Experimental Results
COMIC (Ours) $\mathbf{1.00}$ $\mathbf{1.00}$ $\mathbf{1.00}$ $\mathbf{1.00}$ $\mathbf{1.00}$ $\mathbf{0.97}$ $\mathbf{0.99}$ $\mathbf{1.00}$ $\mathbf{1.00}$ $\mathbf{0.98}$ $\mathbf{1.00}$ $\mathbf{0.97}$ $\mathbf{0.97}$
Sloppy-AIC $0.84$ $0.27$ $0.80$ $0.03$ $0.00$ $0.60$ $0.68$ $0.58$ $0.67$ $0.97$ $0.78$ $0.62$ $0.79$
Sloppy-BIC $\mathbf{1.00}$ $\mathbf{1.00}$ $\mathbf{1.00}$ $0.58$ $\mathbf{1.00}$ $0.74$ $0.74$ $0.90$ $0.90$ $0.59$ $0.88$ $0.54$ $0.59$
Sloper $0.97$ $0.60$ $0.87$ $0.01$ $0.03$ $0.67$ $0.71$ $0.68$ $0.93$ $0.96$ $0.84$ $0.62$ $0.79$
Slope $0.09$ $0.23$ $0.12$ $0.03$ $0.01$ $0.48$ $0.57$ $0.45$ $0.46$ $0.97$ $0.68$ $0.59$ $0.81$
QCCD $\mathbf{1.00}$ $0.90$ $1.00$ $1.00$ $1.00$ $0.67$ $0.80$ $0.71$ $0.91$ $0.56$ $0.91$ $0.56$ $0.79$
IGCI-Uniform $0.17$ $0.37$ $0.51$ $0.42$ $0.02$ $0.38$ $0.43$ $0.57$ $0.51$ $\mathbf{0.98}$ $0.59$ $0.58$ $0.69$
IGCI-Gaussian $0.98$ $\mathbf{1.00}$ $0.99$ $0.99$ $0.94$ $0.32$ $0.38$ $0.95$ $0.66$ $0.78$ $0.57$ $0.56$ $0.64$
GPLVM $\mathbf{1.00}$ $\mathbf{1.00}$ $\mathbf{1.00}$ $\mathbf{1.00}$ $\mathbf{1.00}$ $0.92$ $0.91$ $0.99$ $0.97$ $\mathbf{0.98}$ $0.99$ $0.82$ $0.80$
LOCI-HSIC $\mathbf{1.00}$ $\mathbf{1.00}$ $0.99$ $0.96$ $\mathbf{1.00}$ $0.86$ $0.90$ $0.89$ $0.82$ $0.81$ $0.95$ $0.74$ $0.49$
LOCI-Likelihood $\mathbf{1.00}$ $\mathbf{1.00}$ $\mathbf{1.00}$ $\mathbf{1.00}$ $\mathbf{1.00}$ $0.53$ $0.56$ $0.85$ $0.91$ $0.88$ $0.82$ $0.45$ $0.65$
CAM $\mathbf{1.00}$ $\mathbf{1.00}$ $\mathbf{1.00}$ $0.35$ $0.82$ $0.62$ $0.65$ $0.88$ $0.86$ $0.38$ $0.82$ $0.59$ $0.38$
RESIT $\mathbf{1.00}$ $\mathbf{1.00}$ $0.65$ $0.03$ $0.00$ $0.78$ $0.81$ $0.75$ $0.80$ $0.55$ $0.79$ $0.82$ $0.68$
@@ Appendix G Ablation Studies
In this part, we conduct additional ablation studies to evaluate the contribution of the location-scale estimation and the additional optimization of the model complexity to the performance of our COMIC approach.
Location-Scale Estimation and Location-Only Estimation
In this part of the ablation studies, we compare the location-scale estimation to the location-only estimation for the Gaussian likelihood. Instead of modeling the location and scale with the neural network, we only learn the location and fix the scale value to $1.0$. The results of the two approaches are illustrated in Fig. 2. On easier synthetic benchmarks from Tagasovska et al. (2020), comprising AN, AN-s, LS, LS-s, and MN-U, there is no difference in performance as both approaches attain the ideal results on every set. However, as the datasets become more complicated, the gaps of performance between them are more noticeable. On the SIM-c set with hidden confounders, the increase in fitness of the location-scale modeling can allow for better predictions. In particular, on the CE-Multi and Tübingen benchmarks, the improvements in performance are more significant. On CE-Multi, due to the diversity in underlying causal models, the flexibility of the location-scale COMIC can be advantageous. With the more realistic causal relations in the Tübingen cause-effect pairs, the increase in fitness also yields better results with higher confidences, which are evidenced with close to ideal Bi-AUROC. Hence, although the location-scale estimation introduces higher model complexity, the enhancement in fitness is more beneficial in this case.
Figure 2: Performance of COMIC with location-scale (solid lines) and location-only (dashed lines) models across all benchmarks. Higher accuracy and bidirectional AUROC (Bi-AUROC) are preferable. Despite higher model complexity, the location-scale modeling approach enables more accurate predictions of causal relations on more complicated datasets.
Importance of Model Complexity
Since in the previous ablation study, the increase of fitness can improve the overall performance, in this part, we study the importance of considering the model complexity in determining the causal direction. We compare the same network architecture and training hyperparameters with two different objectives, considering both model fitness and complexity and only considering the model fitness, whose results are depicted in Fig. 3. Analogous to the previous ablation study, the scores on the first five benchmarks (from AN to MN-U) are comparable. Despite that, the distinctions are more pronounced on the remaining datasets. When the model complexity is not considered in optimization, the performance is substantially decreased. Moreover, the greater the divergence between the ground truth causal relations and the assumed models, such as those in SIM, SIM-c, CE-Cha, and Tübingen, the larger the performance gap between them will become. From this study, it is obvious that the variational Bayesian approach to optimizing the parameters is more effective than the point estimation one which only tries to minimize the negative log-likelihoods.
Figure 3: Performance of COMIC with (solid lines) and without (dotted lines) considering the model complexity across all benchmarks. Higher accuracy and bidirectional AUROC (Bi-AUROC) are preferable. Considering the model complexity in addition to model fitness when learning the conditional distributions contributes significantly to the ability of our approach to accurately distinguish the causal directions.
@@
Coker et al. (2022) have indicated that mean-field variational inference (MF-VI) of wide Bayesian neural networks can cause the network to ignore the data, where the variational posteriors will collapse to the priors instead of the true posteriors. In this part, we examine the behavior of our models with different hidden layer widths. The result is depicted in Tab. 2. From these results, there are impacts on performance when the width of the hidden layer increases. When the width reaches $100$ and $200$, the reductions in accuracy scores become noticeable. Nonetheless, the Bi-AUROC scores still remain satisfactory. At $500$ hidden nodes, we begin to experience severer declines in performance. However, it is important to note that such excessively large widths are not recommended in the bivariate setting, where there is only one input dimension and up to two output dimensions. Thus, our approach will not be substantially affected by the ignorance of the data (Coker et al., 2022), if the width of the hidden layer stays within an appropriate range ($\leq 50$).
Table 2: The effect of hidden layer width on the performance of COMIC. Higher accuracy and bidirectional AUROC (Bi-AUROC) are preferable. The impact of the width on the Bi-AUROC is tolerable up until $200$ nodes. When the width reaches $500$—a discouraged choice in the bivariate setting—the decline in performance is severe.
@@
[…]
$100$ $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $0.89$ $0.90$ $0.97$ $1.00$ $0.89$ $0.95$ $0.87$ $0.78$
$200$ $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $0.88$ $0.90$ $0.97$ $1.00$ $0.88$ $0.95$ $0.87$ $0.78$
$500$ $1.00$ $1.00$ $0.92$ $1.00$ $1.00$ $0.83$ $0.74$ $0.90$ $0.82$ $0.82$ $0.89$ $0.83$ $0.64$
[…]
$100$ $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $0.97$ $0.97$ $0.99$ $1.00$ $0.97$ $0.99$ $0.95$ $0.91$
$200$ $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $0.97$ $0.97$ $0.99$ $1.00$ $0.97$ $0.99$ $0.95$ $0.91$
$500$ $1.00$ $1.00$ $0.99$ $1.00$ $1.00$ $0.93$ $0.85$ $0.97$ $0.93$ $0.93$ $0.96$ $0.93$ $0.77$
Number of Layers of Bayesian Neural Networks
We have also studied the effect of adding another hidden layer for the neural networks with the results presented in Tab. 3. With a two-layer configuration with $10,5$ hidden nodes, the neural networks have the same number of parameters as the ones with one hidden layer of $20$ nodes. The obtained results show no substantial distinctions in performance when an additional hidden layer is involved. Additionally, the Bi-AUROC scores on some datasets, such as SIM, SIM-c, and Tübingen, are slightly decreased. The negligible changes in performance indicate that neural networks with one hidden layer is adequate for the bivariate causal discovery task, and the inclusion of additional layers should be unnecessary.
Table 3: The effect of adding another hidden layer on the performance of COMIC. Higher accuracy and bidirectional AUROC (Bi-AUROC) are preferable. The neural networks with $10,5$ hidden nodes have the same number of parameters as the ones with $20$ hidden nodes. There are no significant differences in performance when one more layer is included in the neural networks.
Hidden Nodes AN AN-s LS LS-s MN-U SIM SIM-c SIM-G SIM-ln CE-Multi CE-Net CE-Cha Tübingen
[…]
$20$, $00$ $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $0.90$ $0.93$ $0.99$ $1.00$ $0.89$ $0.97$ $0.90$ $0.87$
[…]
$10$, $10$ $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $0.89$ $0.93$ $0.98$ $1.00$ $0.90$ $0.96$ $0.91$ $0.87$
[…]
$20$, $10$ $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $0.87$ $0.91$ $0.99$ $1.00$ $0.91$ $0.97$ $0.90$ $0.89$
[…]
$20$, $00$ $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $0.97$ $0.99$ $1.00$ $1.00$ $0.98$ $1.00$ $0.97$ $0.97$
[…]
$10$, $10$ $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $0.97$ $0.99$ $1.00$ $1.00$ $0.98$ $1.00$ $0.98$ $0.96$
[…]
$20$, $10$ $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $0.96$ $0.98$ $1.00$ $1.00$ $0.98$ $1.00$ $0.97$ $0.96$
@@
Besides the MAP approach for selecting the hyperparameters in App. B, the sparsity-inducing prior proposed by Louizos et al. (2017) can be considered as an alternative option for the hyperprior. In this prior, the scale hyperparameters follow the non-informative Jeffreys hyperprior $p\left(z\right)\propto\left|z\right|^{-1}$ instead of a uniform hyperprior. With this choice of prior, the posteriors of the hyperparameters are inferred using MF-VI with Gaussian distributions adopted as the variational posteriors (Louizos et al., 2017). The Gaussian posteriors of the hyperparameters enable a pruning mechanism where the weights with $\log\sigma_{z}^{2}-\log\mu_{z}^{2}$ exceeding a threshold $t$ are pruned from the network. We compare the results between these two hyperpriors in Tab. 4. The results indicate that there is no significant difference in result exhibited on synthetic datasets. On the Tübingen benchmark, slight divergences are observed in the accuracy scores. However, the variations in Bi-AUROC scores remain negligible.
Table 4: Performance of COMIC with uniform and Jeffreys hyperpriors (with and without the post-hoc pruning, Louizos et al., 2017). Higher accuracy and bidirectional AUROC (Bi-AUROC) are preferable. There are no substantial differences in accuracy and Bi-AUROC scores, except for the Tübingen benchmark.
@@
Uniform $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $0.90$ $0.93$ $0.99$ $1.00$ $0.89$ $0.97$ $0.90$ $0.87$
Jeffreys w/o Pruning $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $0.90$ $0.93$ $0.98$ $1.00$ $0.90$ $0.97$ $0.90$ $0.84$
Jeffreys w/ Pruning $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $0.89$ $0.93$ $0.98$ $1.00$ $0.90$ $0.96$ $0.89$ $0.91$
@@
Uniform $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $0.97$ $0.99$ $1.00$ $1.00$ $0.98$ $1.00$ $0.97$ $0.97$
Jeffreys w/o Pruning $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $0.97$ $0.99$ $1.00$ $1.00$ $0.97$ $1.00$ $0.97$ $0.96$
Jeffreys w/ Pruning $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $0.97$ $0.99$ $1.00$ $1.00$ $0.97$ $1.00$ $0.97$ $0.97$
@@
While the standard Gaussian distribution is a common choice in previous methods for encoding the marginal distributions of the causes, as discussed in Sec. 4.3, we also explore alternative approaches. Specifically, we compare our choice against other models for computing codelengths including the uniform prior based on the data range, the uniform prior based on the data resolution as proposed by Marx & Vreeken (2017, 2019b), and the variational Bayesian Gaussian mixture model (VB-GMM, Blei & Jordan, 2006) with a maximum of $50$ components as in Mooij et al. (2010). As shown in the results in Tab. 5, using uniform codelengths generally degrades performance across most datasets. While VB-GMM codelengths offer improvements on the SIM, CE-Multi, and CE-Net synthetic benchmarks, they also lead to noticeable performance drops on SIM-G, SIM-ln, and CE-Cha. For the real-world Tübingen dataset, although VB-GMM yields a notable increase in prediction accuracy, it also results in a substantial decrease in Bi-AUROC. Overall, these results support the use of the standard Gaussian as a reasonable choice for our method. Furthermore, adopting more flexible models for encoding the marginals could also hinder the theoretical analysis of separable-compatibility. However, using the standard Gaussian codelength can introduce an inductive bias toward “more Gaussian” causes, which should be carefully considered in more complex scenarios, such as multivariate settings or those involving hidden confounders.
Table 5: The effect of different cause encoding choices on the performance of COMIC. Higher accuracy and bidirectional AUROC (Bi-AUROC) are preferable. Our usage of the standard Gaussian codelength is compared against several alternative encoding methods for the causes: the uniform prior based on the data range, the uniform prior based on the data resolution (Marx & Vreeken, 2017, 2019b), and the variational Bayesian Gaussian mixture model (VB-GMM, Blei & Jordan, 2006). The uniform and VB-GMM codelengths introduce varied performance changes, with VB-GMM showing improved performance on some synthetic benchmarks but reduced performance on others, and notably decreasing Bi-AUROC on the Tübingen dataset despite higher accuracy.
@@
Standard Gaussian $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $0.90$ $0.93$ $0.99$ $1.00$ $0.89$ $0.97$ $0.90$ $0.87$
Uniform (Range) $1.00$ $0.93$ $1.00$ $1.00$ $1.00$ $0.88$ $0.84$ $0.83$ $0.99$ $0.94$ $0.94$ $0.80$ $0.86$
Uniform (Resolution) $0.94$ $0.75$ $0.97$ $0.95$ $0.79$ $0.59$ $0.68$ $0.64$ $0.98$ $0.86$ $0.75$ $0.79$ $0.80$
VB-GMM $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $0.98$ $0.94$ $0.93$ $1.00$ $0.97$ $1.00$ $0.85$ $0.91$
@@
Standard Gaussian $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $0.97$ $0.99$ $1.00$ $1.00$ $0.98$ $1.00$ $0.97$ $0.97$
Uniform (Range) $1.00$ $0.98$ $1.00$ $1.00$ $1.00$ $0.95$ $0.91$ $0.93$ $1.00$ $0.99$ $0.98$ $0.89$ $0.95$
Uniform (Resolution) $0.99$ $0.81$ $0.99$ $1.00$ $0.87$ $0.66$ $0.74$ $0.75$ $1.00$ $0.95$ $0.84$ $0.88$ $0.81$
VB-GMM $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $0.98$ $0.98$ $1.00$ $1.00$ $1.00$ $0.90$ $0.90$

### v4

@@ Appendix F Detailed Experimental Results
COMIC (Ours) $\mathbf{1.00}$ $\mathbf{1.00}$ $\mathbf{1.00}$ $\mathbf{1.00}$ $\mathbf{1.00}$ $0.58$ $0.59$ $0.85$ $0.95$ $0.91$ $0.84$ $0.47$ $0.75$ $\mathit{0.843\pm 0.186}$
Sloppy-AIC $0.84$ $0.27$ $0.80$ $0.03$ $0.00$ $0.60$ $0.68$ $0.58$ $0.67$ $0.97$ $0.78$ $0.62$ $0.79$ $0.587\pm 0.301$
Sloppy-BIC $\mathbf{1.00}$ $\mathbf{1.00}$ $\mathbf{1.00}$ $0.58$ $\mathbf{1.00}$ $0.74$ $0.74$ $0.90$ $0.90$ $0.59$ $0.88$ $0.54$ $0.59$ $0.805\pm 0.182$
Sloper $0.97$ $0.60$ $0.87$ $0.01$ $0.03$ $0.67$ $0.71$ $0.68$ $0.93$ $0.96$ $0.84$ $0.62$ $0.79$ $0.668\pm 0.313$
Slope $0.09$ $0.23$ $0.12$ $0.03$ $0.01$ $0.48$ $0.57$ $0.45$ $0.46$ $0.97$ $0.68$ $0.59$ $0.81$ $0.422\pm 0.307$
QCCD $\mathbf{1.00}$ $0.90$ $1.00$ $1.00$ $1.00$ $0.67$ $0.80$ $0.71$ $0.91$ $0.56$ $0.91$ $0.56$ $0.79$ $0.833\pm 0.164$
IGCI-Uniform $0.17$ $0.37$ $0.51$ $0.42$ $0.02$ $0.38$ $0.43$ $0.57$ $0.51$ $\mathbf{0.98}$ $0.59$ $0.58$ $0.69$ $0.478\pm 0.235$
IGCI-Gaussian $0.98$ $\mathbf{1.00}$ $0.99$ $0.99$ $0.94$ $0.32$ $0.38$ $0.95$ $0.66$ $0.78$ $0.57$ $0.56$ $0.64$ $0.750\pm 0.244$
GPLVM $\mathbf{1.00}$ $\mathbf{1.00}$ $\mathbf{1.00}$ $\mathbf{1.00}$ $\mathbf{1.00}$ $\mathbf{0.92}$ $\mathbf{0.91}$ $\mathbf{0.99}$ $\mathbf{0.97}$ $\mathbf{0.98}$ $\mathbf{0.99}$ $\mathbf{0.82}$ $\mathbf{0.80}$ $\mathbf{0.951\pm 0.069}$
LOCI $\mathbf{1.00}$ $\mathbf{1.00}$ $\mathbf{1.00}$ $\mathbf{1.00}$ $\mathbf{1.00}$ $0.53$ $0.56$ $0.85$ $0.91$ $0.88$ $0.82$ $0.45$ $0.65$ $0.819\pm 0.202$
CAM $\mathbf{1.00}$ $\mathbf{1.00}$ $\mathbf{1.00}$ $0.35$ $0.82$ $0.62$ $0.65$ $0.88$ $0.86$ $0.38$ $0.82$ $0.59$ $0.38$ $0.719\pm 0.240$
@@ Appendix G Ablation Studies
Coker et al. (2022) have indicated that mean-field variational inference (MF-VI) of wide Bayesian neural networks can cause the network to ignore the data, where the variational posteriors will collapse to the priors instead of the true posteriors. In this part, we examine the behavior of our models with different hidden layer widths. The result is depicted in Tab. 2. From these results, there are impacts on performance when the width of the hidden layer increases. When the width reaches $100$, the reductions in accuracy scores become noticeable. Nonetheless, the Bi-AUROC scores still remain satisfactory. At $200$ hidden nodes, we begin to experience severer declines in performance. However, it is important to note that such excessively large width is not recommended in the bivariate setting, where there is only one input dimension and up to two output dimensions. Thus, our approach will not be substantially affected by the ignorance of the data (Coker et al., 2022), if the width of the hidden layer stays within an appropriate range ($\leq 100$).
Table 2: The effect of hidden layer width on the performance of COMIC. Higher accuracy and bidirectional AUROC (Bi-AUROC) are preferable. The impact of the width on the Bi-AUROC is tolerable up until $100$ nodes. When the width reaches 200200—a discouraged choice in the bivariate setting—the decline in performance is severe.
@@
$\phantom{0}10$ $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $0.58$ $0.58$ $0.90$ $0.92$ $0.90$ $0.83$ $0.46$ $0.53$
$\phantom{0}20$ $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $0.58$ $0.59$ $0.89$ $0.96$ $0.90$ $0.83$ $0.46$ $0.68$
$\phantom{0}50$ $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $0.58$ $0.59$ $0.85$ $0.95$ $0.91$ $0.84$ $0.47$ $0.75$
$100$ $1.00$ $0.99$ $0.99$ $1.00$ $0.99$ $0.62$ $0.60$ $0.84$ $0.95$ $0.93$ $0.80$ $0.49$ $0.70$
$200$ $0.99$ $0.97$ $0.97$ $0.99$ $0.96$ $0.62$ $0.57$ $0.81$ $0.93$ $0.92$ $0.76$ $0.48$ $0.69$
@@
In addition to the MAP approach for selecting the hyperparameters in App. B, the sparsity-inducing prior proposed by Louizos et al. (2017) can be used as an alternative hyperprior. In this prior, the scale hyperparameters follow the non-informative Jeffreys hyperprior $p\left(z\right)\propto\left|z\right|^{-1}$ rather than a uniform hyperprior. With this choice of prior, the posteriors of the hyperparameters are inferred using MF-VI with Gaussian distributions adopted as the variational posteriors (Louizos et al., 2017). The Gaussian posteriors of the hyperparameters enable a pruning mechanism in which the weights satisfying $\log\sigma_{z}^{2}-\log\mu_{z}^{2}>t$ are pruned from the network. We compare the results between these two hyperpriors in Tab. 3. The results indicate that the uniform hyperprior is generally more effective in comparison to the sparsity-inducing Jeffreys hyperprior on most benchmarks, except for SIM-c. Moreover, pruning yields no significant differences in scores compared to the unpruned case of the Jeffreys hyperprior.
Table 3: Performance of COMIC with uniform and Jeffreys hyperpriors (with and without the post-hoc pruning, Louizos et al., 2017). Higher accuracy and bidirectional AUROC (Bi-AUROC) are preferable. The uniform hyperpriors provide better performance on most datasets, except for SIM-c.
@@
Uniform $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $0.57$ $0.54$ $0.78$ $0.89$ $0.79$ $0.78$ $0.43$ $0.67$
Jeffreys w/o Pruning $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $0.54$ $0.62$ $0.66$ $0.85$ $0.84$ $0.70$ $0.47$ $0.58$
Jeffreys w/ Pruning $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $0.55$ $0.62$ $0.67$ $0.85$ $0.84$ $0.70$ $0.47$ $0.58$
@@
Uniform $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $0.58$ $0.59$ $0.85$ $0.95$ $0.91$ $0.84$ $0.47$ $0.75$
Jeffreys w/o Pruning $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $0.58$ $0.63$ $0.80$ $0.93$ $0.94$ $0.80$ $0.47$ $0.63$
Jeffreys w/ Pruning $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $0.57$ $0.62$ $0.80$ $0.92$ $0.94$ $0.80$ $0.47$ $0.63$
@@
While the standard Gaussian distribution is a common choice in previous methods for encoding the marginal distributions of the causes, as discussed in Sec. 4.3, we also investigate a more advanced alternative for modeling the marginal distribution of the causes, which is the variational Bayesian Gaussian mixture model (VB-GMM, Blei & Jordan, 2006). The VB-GMM is a flexible model that can capture multimodal distributions, which may be beneficial in scenarios where the causes exhibit complex distributions. Despite that, the identifiability of the causal direction is more challenging to evaluate theoretically with the VB-GMM. As shown in Tab. 4, VB-GMM substantially improves prediction accuracy and Bi-AUROC over the standard Gaussian on certain synthetic benchmarks, such as SIM, SIM-c, and CE-Net. This improvement, however, is not uniform across datasets: on SIM-G and CE-Multi, performance drops notably, while on the real-world Tübingen dataset, VB-GMM increases accuracy but slightly reduces Bi-AUROC.
Table 4: The effect of different cause encoding choices on the performance of COMIC, including the standard Gaussian and variational Bayesian Gaussian mixture models (VB-GMM) encoding methods. Higher accuracy and bidirectional AUROC (Bi-AUROC) are preferable. Using the more advanced VB-GMM method for encoding the causes yields mixed results, showing significant improvements in performance on some datasets, but reducing on some others, including the real-world Tübingen benchmark.
@@
Standard Gaussian $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $0.57$ $0.54$ $0.78$ $0.89$ $0.79$ $0.78$ $0.43$ $0.67$
VB-GMM $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $0.77$ $0.71$ $0.57$ $0.90$ $0.74$ $0.86$ $0.48$ $0.73$
@@
Standard Gaussian $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $0.58$ $0.59$ $0.85$ $0.95$ $0.91$ $0.84$ $0.47$ $0.75$
VB-GMM $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $0.80$ $0.79$ $0.67$ $0.95$ $0.81$ $0.96$ $0.47$ $0.73$